The perimeter of rectangle is 42 m. The length of the rectangle is 3 m more than twice the width. Find the dimensions of the rectangle.
Answer by ewatrrr(24785)


Hi
The length of the rectangle is 3m more than twice the width
Let x and (2x+3m)represent the width and length respectively
Question states***
2(2x+ 3m) + 2x = 42m
Solving for x
6x = 36m
x = 6m, the width. the length is 15m. (2*6m + 3m)
CHECKING our Answer***
2*15m + 2*6m = 30m + 12m = 42cm